Direct Materials Variances The following data relate to the direct materials cost for the production of 2,500 automobile tires: Actual: 53,500 lbs. at $1.85 per lb. Standard: 55,100 lbs. at $1.8 per lb. a. Determine the direct materials price variance, direct materials quantity variance, and total direct materials cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number. Direct Materials Price Variance $ 2,675 Unfavorable Direct Materials Quantity Variance $ Favorable Total Direct Materials Cost Variance $ Favorable
Answer #1

Direct material price variance = Actual quantity*Standard price - Actual quantity*Actual price

Direct material price variance = 53,500*$1.8 - 53,500*$1.85

Direct material price variance = $96,300 - 98,975 = $2,675 Unfavorable

Direct material quantity variance = Standard quantity*Standard price - Actual quantity*Standard price

Direct material quantity variance = 55,100*$1.8 - 53,500*$1.8

Direct material quantity variance = $99,180 - 96,300 = $2,880 Favorable

Total direct material cost variances = Standard quantity*Standard price - Actual quantity*Actual price

Total direct material cost variances = $99,180 - 98,975 = $205 Favorable